# Channel Knowledge Map-assisted Dual-domain Tracking and Predictive Beamforming for High-Mobility Wireless Networks

### v1

@@ Abstract
Keywords:
Tracking, dual-domain, channel knowledge map, extended Kalman filter, high-mobility, predictive beamforming.
@@ I-A Background
In high-mobility wireless networks, predictive BF is critical to maintaining consistent link quality, as the procedures for CSI acquisition and BF design require substantial training overhead. In particular, based on predicted CSI or multipath AoA information, the transmitter designs BF strategies preemptively that optimize system communication and/or sensing performance for upcoming time slots. However, most existing works on predictive BF algorithms assume the presence of a LoS path [15], a condition that often dues not hold in complex environments, such as urban vehicular networks. In these settings, LoS path might be obstructed by tall buildings or other vehicles and NLoS paths are typicality inevitable. Moreover, as analyzed in [7], BF directly towards the AoA of the LoS path is not always the optimal strategy in terms of achievable rate, with a performance gap of nearly one-third compared to optimal BF in a multipath channel. Indeed, effective predictive BF design in NLoS propagation environments enables channel utilization and provides a stronger received signal/echo power, which in turn improves the beam tracking performance. Building on this insight, this paper further designs a predictive BF strategy that provides high-quality measurements for the proposed CKM-assisted dual-domain tracking scheme.
@@ I-C Our Contributions
  • We propose a CKM-assisted dual-domain tracking framework that effectively integrates information from both the coordinate and beam domains. In the C-Domain, the location and velocity of the target are the states to be updated, with the estimated channel parameters from the B-Domain acting as measurements, directly for LoS-present cases and indirectly via CKM for LoS-absent cases. In the B-Domain, the received echo serves as the measurement and the multipath AoAs are the state to be updated, via combining the a priori information provided by the C-Domain through CKM. This integration can enhance beam tracking accuracy by exploiting CKM to reference the tracked state of the moving target and improve target tracking accuracy by inversely leveraging CKM with the tracked CSI.
  • In the C-Domain, we propose an EKF-based target tracking algorithm to predict and track the state of a moving user terminal, with state evolution and measurement models defined for both LoS-present and LoS-absent conditions. Even when the LoS path is obscured, the proposed approach can maintain continuous target tracking by exploiting NLoS path information through the CKM. In the B-Domain, depending on the condition of the beam transition, we propose an adaptive a priori information fusion strategy for creating the beam transfer probability matrices (TPMs) to assist beam tracking, which effectively combines the channel temporal correlation and the channel a priori information provided by CKM. The proposed approach can track the fast and even burst changes of multipath AoA due to the dynamic environment by exploiting CKM.
@@
  • We conduct extensive simulations to evaluate the superior performance of the proposed dual-domain tracking and predictive BF scheme. Our simulation results demonstrate that the proposed scheme achieves both accurate and continuous target and beam tacking even in complex high-mobility wireless networks, while traditional tracking schemes often fail. Besides, the proposed predictive BF and power allocation design further improves the beam tracking accuracy, particularly for NLoS paths.
@@ II System Model
Consider a downlink high-mobility wireless vehicular network [^1] Note that the considered vehicular network is a typical high-mobility network and the proposed dual-domain tracking framework can be extended to other high-mobility wireless networks, such as UAV networks. as depicted in Fig. 1. A roadside unit (RSU) serves as both a communication transmitter and a sensing receiver, providing services to a moving vehicle. The full-duplex operation RSU is equipped with $N_{t}$ transmitting antennas and $N_{r}$ receiving antennas, both constituting uniform linear arrays (ULAs) with half wavelength antenna spacing. The vehicle is equipped with a ULA of $N_{s}$ antenna elements for receiving signals from the RSU. For ease of exposition, as adopted in the literature [18, 9, 5, 7], we assume that the vehicle travels along the road parallel to the antenna array of RSU, initially opposite and subsequently moving away from the RSU [^2] To facilitate the presentation of the proposed dual-domain tracking framework, we consider a simple tracking scenario in this paper. However, it is worth noting that the proposed framework can be extended to accommodate more complicated tracking scenarios, such as turning vehicles.. In the considered system model, the position of the vehicle is represented by $\boldsymbol{q}=[q_{x},q_{y}]$ and its velocity is denoted as $v_{1}$. We assume that the velocity is a scalar, with its sign indicating movement direction along the x-axis, defined as the horizontal axis in Fig. 1.
@@
Assume $T_{\max}$ being the maximum time duration of interest. The time period $T_{\max}$ is divided into several time slots of length $\Delta T$, with each slot further divided into $L$ symbol periods. Within each slot, we assume that the channel parameters remain unchanged [18, 5, 30]. The transmitted $N_{s}$ data streams from the RSU in a slot are given by $\boldsymbol{S}=[\boldsymbol{s}^{T}_{1};\cdots;\boldsymbol{s}^{T}_{N_{s}}]=[\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{L}]\in\mathbb{C}^{N_{s}\times L}$, where $\boldsymbol{s}_{n_{s}}\in\mathbb{C}^{L\times 1}$, ${n_{s}}\in\{1,\ldots,N_{s}\}$, denotes the $n_{s}$-th downlink data streams and $\boldsymbol{x}_{l}\in\mathbb{C}^{N_{s}\times 1}$, denotes the transmitted signal at the $l$-th snapshot, $\forall l\in\{1,\ldots,L\}$. We assume $\boldsymbol{S}_{i,j}\sim\mathcal{CN}(0,1)$, $\forall i,j$, are independent and identically distributed (i.i.d.), leading to $\mathbb{E}[\boldsymbol{x}_{i}\boldsymbol{x}_{j}^{H}]=\delta(i-j)\boldsymbol{I}_{N_{s}}$, $\forall i,j$. Owing to the independence among different data streams, for sufficiently large $L$ we have $\boldsymbol{s}^{H}_{n_{s}}\boldsymbol{s}_{n_{s}^{\prime}}\approx\delta(n_{s}-n_{s}^{\prime})L$. The $N_{s}$ downlink data streams transmitted at the $[\eta]$-th slot are denoted as $\boldsymbol{S}[\eta]=[\boldsymbol{x}_{1}[\eta],\cdots,\boldsymbol{x}_{L}[\eta]]\in\mathbb{C}^{N_{s}\times L}$. The precoded signal in the $[\eta]$-th slot at the RSU is $\tilde{\boldsymbol{S}}[\eta]=\boldsymbol{F}[\eta|\eta-1]\boldsymbol{S}[\eta]\in\mathbb{C}^{N_{t}\times L}$, where $\boldsymbol{F}[\eta|\eta-1]\in\mathbb{C}^{N_{t}\times N_{s}}$ is the predictive BF matrix. In high-mobility scenario, obtaining perfect CSI at the transmitter is challenging. Thus, the predictive BF matrix $\boldsymbol{F}[\eta|\eta-1]$ is designed based on the predicted CSI in the proposed dual-domain tracking framework.
@@
where $i\in\{1,\dots,P\}$ is the path index [^3] For the LoS-present case, path 1 is a LoS path, while path 2 is an NLoS path. For LoS-absent case, path 1 becomes the strongest NLoS path., $\beta_{i}[\eta]\in\mathbb{C}$, $\tau_{i}[\eta]\in\mathbb{R}$ and $\mu_{i}[\eta]\in\mathbb{R}$ denote the channel path gain, the round-trip delay, and the round-trip Doppler frequency shift of the $i$-th path in the $[\eta]$-th slot, respectively. $\boldsymbol{a}(\theta[\eta])$ and $\boldsymbol{b}(\theta[\eta])$ denote the steering vectors of the RSU transmitting and receiving ULAs, respectively, and they are defined by $\boldsymbol{a}(\theta[\eta])=[1,e^{j\pi\cos(\theta[\eta])},\cdots,e^{j\pi(N_{t}-1)\cos(\theta[\eta])}]^{T}\in\mathbb{C}^{N_{t}\times 1}$, and $\boldsymbol{b}(\theta[\eta])=[1,e^{j\pi\cos(\theta[\eta])},\cdots,e^{j\pi(N_{r}-1)\cos(\theta[\eta])}]^{T}\in\mathbb{C}^{N_{r}\times 1}$. The channel path gains $\beta_{i}[\eta]$ captures both effects of the radar cross-section (RCS) of the vehicle/scatterers and the propagation loss. Denote $\theta_{i}[\eta]$ as the angle of the $i$-th path relative to the RSU and we assume that the AoA and angle of departure (AoD) of the $i$-th path at the RSU in the $[\eta]$-th slot are identical, as commonly adopted in [30, 18].
@@
CKM aims to provide location-specific channel parameter tuples for an area of interest. Specifically, CKM is essentially a mapping function from the user terminal location $\boldsymbol{q}$ to the location-specific channel knowledge such as path gain, angles, delay, and Doppler shift444 In the CKM literature [12], the channel Doppler shift can be obtained from a CKM based on the location and velocity information. In practice, the Doppler information can be sampled in a dynamic environment or inferred based on the sampled Doppler shift in a static scattering environment. This paper focused on tracking algorithm design and thus the CKM construction is out of scope. Note that the proposed dual-domain tracking scheme can be straightforwardly extended to the case without Doppler information in the CKM.[12], i.e.,
@@
where $\alpha_{i}\in\mathbb{R}$ captures the propagation power loss of the $i$-th path relative to the RSU. The channel path gain is given by $\beta_{i}=\varepsilon\alpha_{i}^{2}$, and $\varepsilon$ is the time-invariant reflection coefficient of the considered user terminal [6, 18, 4]. To facilitate the presentation in sequel, we assume that the number of paths stored in CKM is the same as the number of paths $P$ in (1). In practice, the number of paths during the time duration of interest might be time-varying and be inconsistent with the number of paths stored in the CKM, requiring further investigation, which is left for future work.
@@ III CKM-assisted Dual-domain Tracking Framework
The proposed framework is well-designed with the traditional Kalman filtering framework [18, 3, 31], which consists of state prediction and update steps across consecutive time slots. Distinctively, unlike existing ones, it is defined in both the beam and coordinate domains, interconnected through the CKM.
@@
In the C-Domain, the measurement models are addressed separately for two scenarios: LoS-present and LoS-absent conditions. Adopt the index $r=1,2$ to denote the LoS-present and LoS-absent cases, respectively. For LoS-present cases, the C-Domain measurement, vector $\boldsymbol{\beta}_{1}[\eta]=[\tau_{1}[\eta],\nu_{1}[\eta],\cos(\theta_{1}[\eta])]^{T}\in\mathbb{C}^{3\times 1}$, comprises the delay, Doppler shift, and the cosine of the AoA for the LoS path. For LoS-absent cases, vector $\boldsymbol{\beta}_{2}[\eta]=[\boldsymbol{\tau}[\eta];\boldsymbol{\mu}[\eta];\cos(\boldsymbol{\theta}[\eta])]\in\mathbb{C}^{3P\times 1}$ is defined by the B-Domain estimated path parameters ${\boldsymbol{\tau}}[\eta]$, ${\boldsymbol{\nu}}[\eta]$, and ${\boldsymbol{\theta}}[\eta]$ for all $P$ paths. On the other hand, the C-Domain states are defined identically for both LoS-present and LoS-absent cases, which consists of the position and velocity of the moving vehicle, i.e., $\boldsymbol{\alpha}[\eta]=\left[\boldsymbol{q}[\eta],v_{1}[\eta]\right]^{T}\in\mathbb{R}^{3\times 1}$. A measurement model exploiting the geometric relationship that connects $\boldsymbol{\alpha}[\eta]$ and $\boldsymbol{\beta}_{1}[\eta]$ is employed in the presence of the LoS path; otherwise, the CKM that connecting multipath parameter estimates and the state of a moving vehicle is utilized to construct the measurement model, as will be detailed in Section IV of this paper. Moreover, an identical C-Domain state evolution model, i.e., a constant velocity movement model, is adopted for both LoS-present and LoS-absent cases. In the $[\eta]$-th time slot, by leveraging the predicted vehicle state $\boldsymbol{\hat{\alpha}}[\eta|\eta-1]$ and the C-Domain measurements $\boldsymbol{\beta}_{r}[\eta],r=(1,2)$, the RSU applies an EKF-based method to update the state of the vehicle, as will be detailed in Section IV of this paper. Then, based on the C-Domain state evolution model, the state of the vehicle is predicted for the $[\eta+1]$-th slot, i.e., $\boldsymbol{\hat{\alpha}}[\eta+1|\eta]$, which is further exploited for state update in the $[\eta+1]$-th time slot.
@@
In the B-Domain, measurement is the received echo signal at the RSU and state corresponds to the multipath AoA. At the beginning of time slot $[\eta]$, the RSU first receives echo signals $\boldsymbol{R}[\eta]$ based on the predictive BF matrix $\boldsymbol{F}[\eta|\eta-1]$, facilitating the extraction of the round trip delay $\boldsymbol{\tau}[\eta]\in\mathbb{R}^{P\times 1}$ and Doppler shift $\boldsymbol{\mu}[\eta]\in\mathbb{R}^{P\times 1}$ through the matched filtering technique [32, 33]. Note that different propagation paths possess different $\tau_{i}[\eta]$ and $\nu_{i}[\eta]$, and thereby the RSU is capable of separating the signals $\boldsymbol{R}_{i}[\eta]$ reflected from different scatterers. For the $i$-th path, the state $\{\theta_{i}[\eta]\}_{i=1}^{P}$ in the B-Domain can be updated by incorporating the state prediction $\{\boldsymbol{p}(\theta_{i}[\eta|\eta-1])\}_{i=1}^{P}$ from the previous slot and the measurement $\boldsymbol{R}_{i}[\eta]$ in the current slot, as detailed in Section V of this paper. The B-Domain state evolution model, which is characterized by the beam TPM, needs to be established from beam tracking. This model relies on the channel’s temporal correlation as well as the a prior information stored in the CKM, where the latter one further depends on the predicted state of the vehicle $\hat{\boldsymbol{\alpha}}[\eta+1|\eta]$. Specifically, in Fig. 2, the matrix $\boldsymbol{\Pi}_{i}^{I}[\eta]$ denotes the TPM for the $i$-th path due to vehicle movement and the CKM generates the beam TPMs $\boldsymbol{\Pi}_{i}^{II}[\eta]$ for each path based on the predicted position $\boldsymbol{\hat{q}}[\eta+1|\eta]$ in the C-Domain. Consequently, the variation in AoA associated with each path can be effectively modeled as a discrete Markov process [6].
@@
Based on the B-Domain state evolution model and the updated B-Domain state $\boldsymbol{\hat{\theta}}[\eta]$, B-Domain state prediction provides two kinds of angle prior information, i.e., the soft prediction $\{\boldsymbol{p}(\theta_{i}[\eta+1|\eta])\}_{i=1}^{P}$ and the hard prediction $\boldsymbol{\hat{\theta}}[\eta+1|\eta]$, where the former is exploited to assist B-Domain state updates while the latter is employed for designing the predictive beamformer. Then, based on the hard prediction $\boldsymbol{\hat{\theta}}[\eta+1|\eta]$, the predictive BF matrix $\boldsymbol{F}[\eta+1|\eta]$ for the next time slot is designed to minimize the maximum AoA estimation error among multiple propagation paths, which determines the received echo signal $\boldsymbol{R}[\eta+1]$ for the next time slot. By incorporating both the soft prediction $\boldsymbol{p}(\theta_{i}[\eta+1|\eta])$ and new measurements $\boldsymbol{R}_{i}[\eta+1]$ in the $[\eta+1]$-th time slot, the B-Domain state $\boldsymbol{\hat{\theta}}[\eta+1]$ can be effectively updated.
@@
Moreover, CKM directly links the coordinate and beam domains and embeds environmental awareness into the tracking process across both domains. From the C-Domain to the B-Domain, CKM enhances beam tracking performance via mapping the state prediction of the vehicle to a prior AoA information, particularly beneficial for NLoS paths characterized by lower sensing signal-to-noise ratios (SNRs). From the B-Domain to the C-Domain, CKM is reversely utilized to infer the state of the vehicle based on multipath parameter measurements for the LoS-absent case. As a result, continuous C-Domain tracking is maintained for both LoS-present and LoS-absent cases. Moreover, since the coherence time of the C-Domain state evolution model significantly exceeds that of the B-Domain state evolution model, C-Domain tracking further improves the continuity in B-Domain tracking via exploiting CKM, particularly scenarios involving random path birth and death, or burst change of multipath AoA.

### v2

@@ I-A Background
In high-mobility wireless networks, predictive BF is critical to maintaining consistent link quality, as the procedures for CSI acquisition and BF design require substantial training overhead. In particular, based on predicted CSI or multipath AoA information, the transmitter designs BF strategies preemptively that optimize system communication and/or sensing performance for upcoming time slots. However, most existing works on predictive BF algorithms assume the presence of a LoS path [31], a condition that often does not hold in complex environments, such as urban vehicular networks. In these settings, LoS path might be obstructed by tall buildings or other vehicles and NLoS paths are typicality inevitable. Moreover, as analyzed in [7], BF directly towards the AoA of the LoS path is not always the optimal strategy in terms of achievable rate, with a performance gap of nearly one-third compared to optimal BF in a multipath channel. Indeed, effective predictive BF design in NLoS propagation environments enables channel utilization and provides a stronger received signal/echo power, which in turn improves the beam tracking performance. Building on this insight, this paper further designs a predictive BF strategy that provides high-quality measurements for the proposed CKM-assisted dual-domain tracking scheme.
@@ I-B Existing Works
Generally, the construction of CKM involves two main stages: data acquisition and map generation [46]. In the data acquisition stage, location-specific channel data can be obtained either from simulation-based methods (e.g., ray tracing over physical environment maps [23]) or from measurement-based approaches, such as offline drive tests and online data collection enabled by the minimization of drive tests (MDT) framework standardized in 3GPP [14]. In the map generation stage, model-free interpolation and model-assisted learning constitute the two main categories of methods for reconstructing [17, 4, 6, 19, 5, 12]. Model-free methods, such as Kriging [17], inverse-distance-weighted (IDW) interpolation [4, 6, 19], and low-rank matrix completion [5] were widely adopted in recent literature as basic techniques for CKM generation. Model-assisted approaches construct the CKM by integrating data-driven learning with geometric models of the propagation environment [12], leveraging ray-tracing-like techniques to infer environmental characteristics with enhanced accuracy and scalability. Compared with more complex Kriging or data-driven learning methods, the IDW method offers an excellent trade-off between the interpolation accuracy and computational scalability, which is essential for large-scale high-mobility simulations.
@@ I-C Our Contributions
  • We propose a CKM-assisted dual-domain tracking framework that effectively integrates information from both the coordinate and beam domains. In the C-Domain, the location and velocity of the target are the states to be updated, with the estimated channel parameters from the B-Domain acting as measurements, directly for LoS-present cases and indirectly via CKM for LoS-absent cases. In the B-Domain, the received echo serves as the measurement and the multipath AoAs are the state to be updated, via combining the a priori information provided by the C-Domain through CKM. This integration can enhance B-Domain tracking accuracy by exploiting CKM to reference the tracked state of the moving target and improve target tracking accuracy by inversely leveraging CKM with the tracked CSI.
  • In the C-Domain, we propose an EKF-based target tracking algorithm to predict and track the state of a moving user terminal, with state evolution and measurement models defined for both LoS-present and LoS-absent conditions. Even when the LoS path is obscured, the proposed approach can maintain continuous target tracking by exploiting NLoS path information through the CKM. In the B-Domain, depending on the condition of the beam transition, we propose an adaptive a priori information fusion strategy for creating the beam transfer probability matrices (TPMs) to assist B-Domain tracking, which effectively combines the channel temporal correlation and the channel a priori information provided by CKM. The proposed approach can track the fast and even burst changes of multipath AoA due to the dynamic environment by exploiting CKM.
@@
  • We conduct extensive simulations to evaluate the superior performance of the proposed dual-domain tracking and predictive BF scheme. Our simulation results demonstrate that the proposed scheme achieves both accurate and continuous target and beam tracking even in complex high-mobility wireless networks, while traditional tracking schemes often fail. Besides, the proposed predictive BF and power allocation design further improves the B-Domain tracking accuracy, particularly for NLoS paths.
@@ II System Model
Consider a downlink high-mobility wireless vehicular network [^1] Note that the considered vehicular network is a typical high-mobility network and the proposed dual-domain tracking framework can be extended to other high-mobility wireless networks, such as UAV networks. as depicted in Fig. 1. A roadside unit (RSU) serves as both a communication transmitter and a sensing receiver, providing services to a moving vehicle. The full-duplex operation RSU is equipped with $N_{t}$ transmitting antennas and $N_{r}$ receiving antennas, both constituting uniform linear arrays (ULAs) with half wavelength antenna spacing. The vehicle is equipped with a ULA of $N_{s}$ antenna elements for receiving signals from the RSU. For ease of exposition, we assume that the vehicle travels along the road parallel to the antenna array of RSU, moving initially towards and subsequently away from the RSU [^2] To facilitate the presentation of the proposed dual-domain tracking framework, we consider a simple tracking scenario in this paper, as adopted in the literature [24, 45, 21, 7]. However, it is worth noting that the proposed framework can be extended to more complicated tracking scenarios. For non-parallel linear movement, the velocity of the vehicle can be represented as a two-dimensional vector with components along the x and y axes, requiring only modifications to the state evolution model and measurement models in the C-Domain. For curved trajectories, the short slot duration allows adopting a constant velocity model between consecutive slots, where the proposed tracking framework is still applicable.. In the considered system model, the position of the vehicle is represented by $\boldsymbol{q}=[q_{x},q_{y}]$ and its velocity is denoted as $v_{1}$. We assume that the velocity is a scalar, with its sign indicating movement direction along the x-axis, defined as the horizontal axis in Fig. 1.
@@
Assume $T_{\max}$ being the maximum time duration of interest. The time period $T_{\max}$ is divided into several time slots of length $\Delta T$, with each slot further divided into $L$ symbol periods. Within each slot, we assume that the channel parameters remain unchanged [24, 21, 44]. The transmitted $N_{s}$ data streams from the RSU in a slot are given by $\boldsymbol{S}=[\boldsymbol{s}^{T}_{1};\cdots;\boldsymbol{s}^{T}_{N_{s}}]=[\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{L}]\in\mathbb{C}^{N_{s}\times L}$, where $\boldsymbol{s}_{n_{s}}\in\mathbb{C}^{L\times 1}$, ${n_{s}}\in\{1,\ldots,N_{s}\}$, denotes the $n_{s}$-th downlink data streams and $\boldsymbol{x}_{l}\in\mathbb{C}^{N_{s}\times 1}$, denotes the transmitted signal at the $l$-th snapshot, $\forall l\in\{1,\ldots,L\}$. We assume $\boldsymbol{S}_{i,j}\sim\mathcal{CN}(0,1)$, $\forall i,j$, are i.i.d., leading to $\mathbb{E}[\boldsymbol{x}_{i}\boldsymbol{x}_{j}^{H}]=\delta(i-j)\boldsymbol{I}_{N_{s}}$, $\forall i,j$. Owing to the independence among different data streams, for sufficiently large $L$ we have $\boldsymbol{s}^{H}_{n_{s}}\boldsymbol{s}_{n_{s}^{\prime}}\approx\delta(n_{s}-n_{s}^{\prime})L$. The $N_{s}$ downlink data streams transmitted at the $\eta]-$th slot are denoted as $\boldsymbol{S}[\eta]=[\boldsymbol{x}_{1}[\eta],\cdots,\boldsymbol{x}_{L}[\eta]]\in\mathbb{C}^{N_{s}\times L}$. The precoded signal in the $\eta]-$th slot at the RSU is $\tilde{\boldsymbol{S}}[\eta]=\boldsymbol{F}[\eta|\eta-1]\boldsymbol{S}[\eta]\in\mathbb{C}^{N_{t}\times L}$, where $\boldsymbol{F}[\eta|\eta-1]\in\mathbb{C}^{N_{t}\times N_{s}}$ is the predictive BF matrix. In high-mobility scenario, obtaining perfect CSI at the transmitter is challenging. Thus, the predictive BF matrix $\boldsymbol{F}[\eta|\eta-1]$ is designed based on the predicted CSI in the proposed dual-domain tracking framework.
@@
where $i\in\{1,\dots,P\}$ is the path index [^3] For the LoS-present case, path 1 is a LoS path, while path 2 is an NLoS path. For LoS-absent case, path 1 becomes the strongest NLoS path., $\beta_{i}[\eta]\in\mathbb{C}$, $\tau_{i}[\eta]\in\mathbb{R}$ and $\mu_{i}[\eta]\in\mathbb{R}$ denote the channel path gain, the round-trip delay, and the round-trip Doppler frequency shift of the $i$-th path in the $\eta]-$th slot, respectively. $\boldsymbol{a}(\theta[\eta])$ and $\boldsymbol{b}(\theta[\eta])$ denote the steering vectors of the RSU transmitting and receiving ULAs, respectively, and they are defined by $\boldsymbol{a}(\theta[\eta])=[1,e^{j\pi\cos(\theta[\eta])},\cdots,e^{j\pi(N_{t}-1)\cos(\theta[\eta])}]^{T}\in\mathbb{C}^{N_{t}\times 1}$, and $\boldsymbol{b}(\theta[\eta])=[1,e^{j\pi\cos(\theta[\eta])},\cdots,e^{j\pi(N_{r}-1)\cos(\theta[\eta])}]^{T}\in\mathbb{C}^{N_{r}\times 1}$. The channel path gains $\beta_{i}[\eta]$ captures both effects of the radar cross-section (RCS) of the vehicle/scatterers and the propagation loss. Denote $\theta_{i}[\eta]$ as the angle of the $i$-th path relative to the RSU and we assume that the AoA and angle of departure (AoD) of the $i$-th path at the RSU in the $\eta]-$th slot are identical, as commonly adopted in [44, 24].
@@
CKM aims to provide location-specific channel parameter tuples for an area of interest. Specifically, CKM is essentially a mapping function from the user terminal location $\boldsymbol{q}$ to the location-specific channel knowledge such as path gain, angles, delay, and Doppler shift [^4] In the CKM literature [46], the channel Doppler shift can be obtained from a CKM based on the location and velocity information. In practice, the Doppler information can be sampled in a dynamic environment or inferred based on the sampled Doppler shift in a static scattering environment. This paper focused on tracking algorithm design and thus the CKM construction is out of scope. Note that the proposed dual-domain tracking scheme can be straightforwardly extended to the case without Doppler information in the CKM.[46], i.e.,
@@
where $\alpha_{i}\in\mathbb{R}$ captures the propagation power loss of the $i$-th path relative to the RSU. The channel path gain is given by $\beta_{i}=\varepsilon\alpha_{i}^{2}$, and $\varepsilon$ is the time-invariant reflection coefficient of the considered user terminal [50, 24, 53]. It is worth noting that, in practice, the CKM is constructed from sampled data using interpolation techniques (as discussed in Section I-B). Consequently, the channel parameters stored in the constructed CKM may deviate slightly from the ground-truth values, thereby introducing CKM construction errors. To facilitate the presentation in sequel, we assume that the number of paths stored in CKM is the same as the number of paths $P$ in (1). In practice, the number of paths during the time duration of interest might be time-varying and be inconsistent with the number of paths stored in the CKM, requiring further investigation, which is left for future work.
@@ III CKM-assisted Dual-domain Tracking Framework
The proposed framework is well-designed with the traditional Kalman filtering framework [24, 36, 38], which consists of state prediction and update steps across consecutive time slots. Distinctively, unlike existing ones, it is defined in both the beam and coordinate domains, interconnected through the CKM. The proposed dual-domain framework aims to establish a CKM-assisted tracking framework that jointly improves both C-Domain and B-Domain tracking accuracy under both LoS-present and LoS-absent conditions. Specifically, the C-Domain tracking aims to maximize the posterior probability of the vehicle state by jointly exploiting the C-Domain prior from the previous slot and the current B-Domain measurements, as will be detailed in Section IV-D. On the other hand, the B-Domain tracking aims to maximize the posterior probability of multipath AoA estimates, utilizing C-Domain state prediction and CKM prior information, as will be detailed in Section V-C. The predictive beamforming is designed to enhance the multipath beam tracking accuracy in the B-Domain, as will be detailed in Section VI, which in turn indirectly improves the target tracking accuracy in the C-Domain through the information exchange assisted by the CKM. The CKM acts as a bridge to enable reciprocal information exchange between the two domains, providing priors for target tracking in the C-Domain via the mapping from multipath channel parameters to potential target positions, and for beam tracking in the B-Domain via the corresponding inverse mapping.
@@
In the C-Domain, the measurement models are addressed separately for two scenarios: LoS-present and LoS-absent conditions. Adopt the index $r=1,2$ to denote the LoS-present and LoS-absent cases, respectively. For LoS-present cases, the C-Domain measurement, vector $\boldsymbol{\beta}_{1}[\eta]=[\tau_{1}[\eta],\nu_{1}[\eta],\cos(\theta_{1}[\eta])]^{T}\in\mathbb{C}^{3\times 1}$, comprises the delay, Doppler shift, and the cosine of the AoA for the LoS path. For LoS-absent cases, vector $\boldsymbol{\beta}_{2}[\eta]=[\boldsymbol{\tau}[\eta];\boldsymbol{\mu}[\eta];\cos(\boldsymbol{\theta}[\eta])]\in\mathbb{C}^{3P\times 1}$ is defined by the B-Domain estimated path parameters ${\boldsymbol{\tau}}[\eta]$, ${\boldsymbol{\nu}}[\eta]$, and ${\boldsymbol{\theta}}[\eta]$ for all $P$ paths. On the other hand, the C-Domain states are defined identically for both LoS-present and LoS-absent cases, which consists of the position and velocity of the moving vehicle, i.e., $\boldsymbol{\alpha}[\eta]=\left[\boldsymbol{q}[\eta],v_{1}[\eta]\right]^{T}\in\mathbb{R}^{3\times 1}$. A measurement model exploiting the geometric relationship that connects $\boldsymbol{\alpha}[\eta]$ and $\boldsymbol{\beta}_{1}[\eta]$ is employed in the presence of the LoS path; otherwise, the CKM that connecting multipath parameter estimates and the state of a moving vehicle is utilized to construct the measurement model, as will be detailed in Section IV of this paper. Moreover, an identical C-Domain state evolution model, i.e., a constant velocity movement model, is adopted for both LoS-present and LoS-absent cases. In the $\eta]-$th time slot, by leveraging the predicted vehicle state $\boldsymbol{\hat{\alpha}}[\eta|\eta-1]$ and the C-Domain measurements $\boldsymbol{\beta}_{r}[\eta],r=(1,2)$, the RSU applies an EKF-based method to update the state of the vehicle, as will be detailed in Section IV of this paper. Then, based on the C-Domain state evolution model, the state of the vehicle is predicted for the $\eta+1]-$th slot, i.e., $\boldsymbol{\hat{\alpha}}[\eta+1|\eta]$, which is further exploited for state update in the $\eta+1]-$th time slot.
@@
In the B-Domain, measurement is the received echo signal at the RSU and state corresponds to the multipath AoA. At the beginning of time slot $\eta]$, the RSU first receives echo signals $\boldsymbol{R}[\eta]$ based on the predictive BF matrix $\boldsymbol{F}[\eta|\eta-1]$, facilitating the extraction of the round trip delay $\boldsymbol{\tau}[\eta]\in\mathbb{R}^{P\times 1}$ and Doppler shift $\boldsymbol{\mu}[\eta]\in\mathbb{R}^{P\times 1}$ through the matched filtering technique [29, 10]. Note that different propagation paths possess different $\tau_{i}[\eta]$ and $\nu_{i}[\eta]$, and thereby the RSU is capable of separating the signals $\boldsymbol{R}_{i}[\eta]$ reflected from different scatterers. For the $i$-th path, the state $\{\theta_{i}[\eta]\}_{i=1}^{P}$ in the B-Domain can be updated by incorporating the state prediction $\{\boldsymbol{p}(\theta_{i}[\eta|\eta-1])\}_{i=1}^{P}$ from the previous slot and the measurement $\boldsymbol{R}_{i}[\eta]$ in the current slot, as detailed in Section V of this paper. The B-Domain state evolution model, which is characterized by the beam TPM, needs to be established for B-Domain tracking. This model relies on the channel’s temporal correlation as well as the a prior information stored in the CKM, where the latter one further depends on the predicted state of the vehicle $\hat{\boldsymbol{\alpha}}[\eta+1|\eta]$. Specifically, in Fig. 2, the matrix $\boldsymbol{\Pi}_{i}^{I}[\eta]$ denotes the TPM for the $i$-th path due to vehicle movement and the CKM generates the beam TPMs $\boldsymbol{\Pi}_{i}^{II}[\eta]$ for each path based on the predicted position $\boldsymbol{\hat{q}}[\eta+1|\eta]$ in the C-Domain. Consequently, the variation in AoA associated with each path can be effectively modeled as a discrete Markov process [50].
@@
Based on the B-Domain state evolution model and the updated B-Domain state $\boldsymbol{\hat{\theta}}[\eta]$, B-Domain state prediction provides two kinds of angle prior information, i.e., the soft prediction $\{\boldsymbol{p}(\theta_{i}[\eta+1|\eta])\}_{i=1}^{P}$ and the hard prediction $\boldsymbol{\hat{\theta}}[\eta+1|\eta]$, where the former is exploited to assist B-Domain state updates while the latter is employed for designing the predictive beamformer. Then, based on the hard prediction $\boldsymbol{\hat{\theta}}[\eta+1|\eta]$, the predictive BF matrix $\boldsymbol{F}[\eta+1|\eta]$ for the next time slot is designed to minimize the maximum AoA estimation error among multiple propagation paths, which determines the received echo signal $\boldsymbol{R}[\eta+1]$ for the next time slot. By incorporating both the soft prediction $\boldsymbol{p}(\theta_{i}[\eta+1|\eta])$ and new measurements $\boldsymbol{R}_{i}[\eta+1]$ in the $\eta+1]-$th time slot, the B-Domain state $\boldsymbol{\hat{\theta}}[\eta+1]$ can be effectively updated.
@@
Moreover, CKM directly links the coordinate and beam domains and embeds environmental awareness into the tracking process across both domains. From the C-Domain to the B-Domain, CKM enhances beam tracking performance via mapping the state prediction of the vehicle to a prior AoA information, particularly beneficial for NLoS paths characterized by lower sensing signal-to-noise ratios (SNRs). From the B-Domain to the C-Domain, CKM is reversely utilized to infer the state of the vehicle based on multipath parameter measurements for the LoS-absent case. As a result, continuous C-Domain tracking is maintained for both LoS-present and LoS-absent cases. Moreover, since the coherence time of the C-Domain state evolution model significantly exceeds that of the B-Domain state evolution model, C-Domain tracking further improves the continuity in B-Domain tracking via exploiting CKM, particularly scenarios involving random path birth and death, or burst change of multipath AoA. However, the uncertainty and inaccuracy of the CKM affect the performance of the proposed dual-domain tracking. In the C-Domain, where the CKM serves as the measurement model for LoS-absent cases, these imperfections directly degrade the accuracy of target state estimation. In the B-Domain, where CKM provides partial prior information for multipath AoA estimation, these imperfections indirectly degrade the accuracy of beam tracking.